OpenStudy (anonymous):

Find the derivative of y=(sin^-2) (x^3)

OpenStudy (shayaan_mustafa):

Hi cmkc109 :) does your question look like\[\large y=\frac{1}{\sin^2(x^3)}\]where x^3 is the angle of sine. right?

OpenStudy (anonymous):

yes.. it is sin ^-2 (x^3) so it is the same as what u wrote there

OpenStudy (anonymous):

ohh..i actually dont know..i thought it is like a product rule or something.. is x^3 suppose to be the angle or sin..

OpenStudy (shayaan_mustafa):

No. it is not a product rule. but a chain rule must be use here. OK.

OpenStudy (anonymous):

ohh..can u still guide me how to do it..still confuse on trig derivatives

OpenStudy (shayaan_mustafa):

sure friend :) \[\large y=\frac{1}{\sin^2(x^3)}\]derivative with respect to x\[\large \frac{d}{dx}(y)=\frac{d}{dx}(\frac{1}{\sin^2(x^3)})\]\[\large \frac{dy}{dx}=\frac{d}{dx}(\sin^{-2}(x^3))\]\[\large \frac{dy}{dx}=(-2)(\cos^-3(x^3))\frac{d}{dx}(x^3)\]\[\large \frac{dy}{dx}=(-2)(\frac{1}{\cos^{3}(x^3)})(3x^2)\]\[\large \frac{dy}{dx}=\frac{-6x^2}{\cos^3(x^3)}\] Here you go buddy :)

OpenStudy (anonymous):

thank you my friend :D

OpenStudy (shayaan_mustafa):

you are ever welcome :)